OpenStudy (awkwardpanda):

Find the values of the variables. [-12, -w^2, 2f, 3] = [2k, -81, -14, 3]

OpenStudy (awkwardpanda):

@jim_thompson5910 im sorry i keep bugging haha. but this one's bugging me

OpenStudy (freckles):

use this: if (x,y)=(a,b) then x=a and y=b.

jimthompson5910 (jim_thompson5910):

freckles has the right idea

jimthompson5910 (jim_thompson5910):

just equate corresponding components

OpenStudy (awkwardpanda):

Thank you (:

jimthompson5910 (jim_thompson5910):

-12 = 2k ----> k = ??? etc etc

OpenStudy (awkwardpanda):

@freckles would -12 = 2k , k=-6?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (awkwardpanda):

YES. Okay so now, the next step would be -w^2 = -81?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (awkwardpanda):

would it be 0.012 or 0.0123?

OpenStudy (ranga):

-w^2 = -81. Multiply by -1: w^2 = 81 w = sqrt(81) = 9

OpenStudy (awkwardpanda):

i was waay off, thank you!

OpenStudy (ranga):

you are welcome.

OpenStudy (awkwardpanda):

and what about 2f = -14. ??

jimthompson5910 (jim_thompson5910):

keep in mind that if w^2 = 81, then w = 9 or w =-9 since those two values make w^2 = 81 true

OpenStudy (freckles):

try to undo that multiplication of 2 by dividing both sides by ?

OpenStudy (awkwardpanda):

right, thanks for that. i fixed it @jim_thompson5910 divide both sides by -14?

jimthompson5910 (jim_thompson5910):